OpenStudy (anonymous):

Find the domain and range of f^-1 given f(x) = cbrt(x+8) + 1

OpenStudy (debbieg):

The inverse function just "swaps" the domain and range of the original function. So if you know the domain and range of \(\Large f(x) =(\sqrt[3]{x+8}) + 1\) Now, that's just a couple of transformations on \(\Large f(x) =\sqrt[3]{x}\). Do you know the domain and range of that? (Or, again, if you know the domain and range of the inverse function \(\Large f(x) =x^3\), then just swap them.

OpenStudy (anonymous):

I think its